# Decoding Quantum LDPC Codes using Collaborative Check Node Removal

### v1

@@ V Conclusion and Future Work
In this work, we propose a collaborative decoding approach for improving the iterative decoding of QLDPC codes through improving qubit separation, with a special focus to GHP codes. We propose a new way to measure the separations of the qubits trapped inside a symmetric stabilizer set. We used this method to go beyond the analysis of qubit separation of the classical trapping sets. Our decoding architecture is free of any post-processing method and can be considered as a two-mode decoder with its main decoder having $\mathcal{O}(n)$ linear complexity. The sub-decoder’s complexity depends on support of the syndrome (an implicit dependence on the physical error rates of the qubits). We discussed for practical situations the complexity of the sub-decoder can be improved as much as almost linear. We showed that inspite of being a post-processing-free decoding the success of the algorithm is a significant improvement compared to the standard iterative min-sum decoder. In future work, we propose on optimizing the current architecture to restore the lost information due to excessive check node removals thereby approaching better logical failure rates.

### v2

@@ V Conclusion and Future Work
In this work, we propose a collaborative decoding approach for improving the iterative decoding of QLDPC codes through improving qubit separation, with a special focus on GHP codes. We propose a new way to measure the separations of the qubits trapped inside a symmetric stabilizer set. We used this method to go beyond the analysis of qubit separation of the classical trapping sets. Our decoding architecture is free of any post-processing method and can be considered as a two-mode decoder with its operation switching the message passing operation of BP over the complete and a modified Tanner graph (or parity check matrix) of the quantum code. The overall decoder’s complexity depends on the support of the syndrome (an implicit dependence on the physical error rate and properties of the code). For GHP codes, it approximately renders a sub-quadratic cost, which, in light of the performance-cost tradeoff, emerges as a practical alternative to the OSD-based post-processing decoders. We showed that, despite being a post-processing-free decoding, the success of the algorithm has a significant improvement compared to the standard iterative min-sum decoder. We would like to conjecture that the QCCNR algorithm does not address the cases of point-like defects. The improved qubit separation can assist only in breaking the trapping set scenarios and is incapable of assisting the iterative decoders when point-like syndrome occurs. Therefore, the slight shift in the decoding performance between QCCNR and BP+OSD00 is due to such point-like syndromes, which might be more reflected in experiments involving small codes like the surface codes. In future work, we will investigate the possible improvements to deal with such point-like defect scenarios and further improve the collaborative architecture’s performance.